OpenStudy (anonymous):

sqrt(x+2)=-2. Solve for x. It looks simple but I just couldn't get it.

sam (.sam.):

sqrt(x+2)=-2 (sqrt(x+2))^(2)=(-2)^(2) x+2=((4)) x+2=(4) x+2=4 x=2

OpenStudy (mertsj):

There is no solution because the equation says that the POSITIVE root is equal to a NEGATIVE number.

OpenStudy (anonymous):

Thats the answer I got a few minutes ago and its not right.

myininaya (myininaya):

no real solution

myininaya (myininaya):

\[\sqrt{f(x)} \ge 0\]

OpenStudy (anonymous):

The answer 2 is what I got.....thanks guys that really helps.

myininaya (myininaya):

never negative

OpenStudy (anonymous):

the answer is certainly not two \[\sqrt{a}\geq0\] so it cannot be -2

OpenStudy (anonymous):

look at the big picture.... think of a something you can take the square root of to get a negative number... if you can't then there is no real solution. there are complex solutions though...

OpenStudy (mertsj):

Just remember that when you solve a radical equation you must always check the answer because the solution process sometimes introduces an extraneous root.

OpenStudy (anonymous):

if you see \[\sqrt{\text{something}}=\text{something negative}\] stop right there

OpenStudy (anonymous):

awesome thanks guys!!! and gals!!